OpenStudy (anonymous):

Find the derivative of : 2cos(2x) + (square root of 3)

OpenStudy (anonymous):

hint the square root of 3 is a number, don't be fooled in to doing other saying "the derivative of a number is zero"

OpenStudy (anonymous):

derivative of cosx = -sinx Try using the Chain rule.

OpenStudy (anonymous):

Okay thank you, I've been told the product rule can help but it's so complicated the way my brother worked it out, I need simpler working out please and possible short cuts. ?

OpenStudy (psymon):

No product rule even in the problem to worry about.

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = 2\frac{ d [ \cos(2x) d(2x)]}{ dx } + 0 \] do that

OpenStudy (anonymous):

Not that complicated the way you wrote it. You want the derivative of \[2\cos(2x) + \sqrt{3}\] or do you want \[2\cos (2x+\sqrt{3})\] or something else?

OpenStudy (anonymous):

The first one :D

OpenStudy (anonymous):

So the first 2 is just a coefficient. What is the derivative of cos?

OpenStudy (anonymous):

Okay , thanks I'll try that first 'mebs', thank you and thanks Paynesdad.

OpenStudy (cggurumanjunath):

-4*sin(2x)